OpenStudy (anonymous):

The corresponding edges of two regular tetrahedrons are 1 cm and 3 cm. If the sum of the weights of the two tetrahedrons is 100 grams and both solids are made up of the same material, find the weight of the bigger solid.

Directrix (directrix):

Any two regular tetrahedrons are similar. The cube of the scale factor is equal to the ratio of the volumes of the two similar solids.

Directrix (directrix):

That means: (1/3)^3 = x/(100 - x) where x is the volume of the smaller solid 1/27 = x / (100 - x)

Directrix (directrix):

Cross multiply to get: 27*x = 100 - x x = ? Check my work

Directrix (directrix):

@saimon0926

Directrix (directrix):

@saimon0926 We are two steps away from an answer here if you will help me a bit.

OpenStudy (anonymous):

thank you :D

Directrix (directrix):

Okay, let's roll. Let me see where we were.

Directrix (directrix):

27*x = 100 - x 28x = 100 7x = 25 x = 25/7 ---> That's x but not the weight of the heavier solid.

Directrix (directrix):

100 - x = 100 - 25/7 = ? If you don't like fractions (I do), you coud solve: 100 - 3.57 = ?

Directrix (directrix):

@saimon0926 Just subtract 3.57 from 100 and that's the answer.

OpenStudy (anonymous):

it would be 96.4 right? :D thnx :D

Directrix (directrix):

Correct!

OpenStudy (anonymous):

thank you so much :)